OpenStudy (anonymous):

is the answer to x^2+7x+10......(x+2)(X+5)?

OpenStudy (anonymous):

yes , it's correct factoring

OpenStudy (katrinakaif):

Coorect =)

OpenStudy (anonymous):

thank you just checking my work

OpenStudy (anonymous):

great and you are welcom. keep on!

OpenStudy (anonymous):

how do i solve x^2-10x+21? i came up with (x-10) (x-1) what did i do wrong?

OpenStudy (anonymous):

(x-7)(x-3)

OpenStudy (anonymous):

First: factor x^2 into ( x ) (x )

OpenStudy (anonymous):

then take the last term i.e +21 , and first check the sign og the middle term

OpenStudy (anonymous):

if its sign minus ==> we got _ and - in both brackets since our last sign +

OpenStudy (anonymous):

and by trail and error method we reach to the right factorization

OpenStudy (anonymous):

and with keeping practicing, it would'nt take much to find it vere easy.

OpenStudy (anonymous):

so is x^2+4x -32.....(x+4) (x-8)?

OpenStudy (anonymous):

it should be (x-4)(x+8) , I'll try to draw it to you too.

OpenStudy (anonymous):

is 2x^2+5xy-12y^2 prime?

OpenStudy (anonymous):



OpenStudy (anonymous):

is 2x^2+5xy-12y^2, what do yo mean by a prime?

OpenStudy (anonymous):

how do i solve this one? isnt 5 prime?

OpenStudy (anonymous):

ok make 5 = 5 *1 or 5 = 1*5

OpenStudy (anonymous):

2x^2+5xy-12y^2 = (x+4y) (2x-3y)

OpenStudy (anonymous):

that is after some trail and error for the numbers and signs till it fits

OpenStudy (anonymous):

i came up with (2x^2+5)(xy-6y^2) as the solution is that correct

OpenStudy (anonymous):

you have to factor the x^2 --- > x*x and the y^2 --->y*y

OpenStudy (anonymous):

oooooo ok

OpenStudy (anonymous):

so when we have here with this little more difficult than the situation before, like what we have 2X^2 , just make 2x*x , wher the 12 y^2 could be many choices ( 4y*3y or 3y*4y , then 2y*6y ot 6y*2y)

OpenStudy (anonymous):

the secret key is that :

OpenStudy (anonymous):

First step : factor the first squared term into two factors of 1st degree

OpenStudy (anonymous):

Second Step: go to the third term and alsoo factor it into 2 multi[liers

OpenStudy (anonymous):

but in step 2 take care , you have to get 2 multipliers that meet the condition of when doing algebracilly sum , the sum or subtract to equal = the middle term with the exact sign

OpenStudy (anonymous):

hope with more practise , no problem you will find, but try to get the skill without any rush

OpenStudy (anonymous):

let the motive inside you mobalize your work. don't give up, it's easy, and with some patcience, things get easier.

OpenStudy (anonymous):

http://www.purplemath.com/modules/solvquad.htm I found this in one of the thread being asked to help.

OpenStudy (anonymous):

thank you!!